# Can femtoscopic correlation function shed light on the nature of the lightest, charm, axial mesons?

### v1

@@ II.2 Scattering lengths and related uncertainties
at a pion mass of 266 MeV. As can be seen, the values for the two channels are very similar. Such a finding can be understood by invoking arguments of heavy quark symmetry. It is important to stress here that a sign convention opposite to that given in Eq. (17) is followed in Ref. Mohler et al. (2013). Further, scattering lengths for systems like $D\bar{K}$, $D_{s}\pi$, $D_{s}K$ and $D\pi$ in isospin 3/2 configurations have been obtained on a lattice in Ref. Liu et al. (2013). Using the values of the low energy constants of the chiral Lagrangian fixed from such a study, the scattering length for $D\pi$ in isospin 1/2 is determined. The results from the former works, obtained at the physical pion mass, can be summarized as
@@
Preliminary results on the $D\pi$ scattering length are available from an alternative source, from the Alice Collaboration Alice ; Grosa , which seem to be in disagreement with the aforementioned values. Let us denote the values given in Refs. Alice ; Grosa as $\tilde{a}^{(I)}$. Using such a notation, the values in Refs. Alice ; Grosa can be summarized as
$\tilde{a}^{(1/2)}_{D\pi}$ $=-0.14\text{ fm to }0.08\text{ fm},$
$\tilde{a}^{(3/2)}_{D^{\pi}}$ $=-0.15\text{ fm to }0.01\text{ fm}.$ (20)
We can expect that $D^{*}\pi$ scattering length would be similar to that for $D\pi$. However, it is not clear which of the two sets of values in Eqs. (19) and (20) should be considered as a benchmark. We find it instructive to investigate both possibilities. In the following discussions, we show that small changes in the value of the mass of the quark-model pole lead to scattering length values compatible with Eq. (19) or (20), though the properties of the two $D_{1}$ states appearing in the amplitudes remain compatible with those of the known states.
@@ II.3 Amplitudes: Results and discussions
As mentioned in the preceding discussions, we add Eq. (16) to the amplitudes already considered in Refs. Gamermann and Oset (2007); Malabarba et al. (2023) with the idea of having the presence of a wide as well as a narrow $D_{1}$ state around 2430 MeV. In addition for the amplitudes to show states with properties compatible with those of $D_{1}(2420)$ and $D_{1}(2430)$, we require the values of the $D^{*}\pi$ scattering lengths to be in agreement with Eq. (19) or (20). For this purpose, we shall consider two different sets of values for the parameters $M_{QM}$ and $g_{QM}$, and refer to the cases as models A and B. It must be clarified here that there is little room for changing the only other possible parameter of the model, which is the subtraction constant used to regularize the loop function, since the properties of $D_{1}(2420)$ are well described by our amplitudes as discussed in Ref. Malabarba et al. (2023). Note that the scale $\mu$ and the subtraction constant $a_{k}$ appearing in Eq. (15) are not two parameters. Both are related to each other and, hence, together they account for one parameter only.
It should be emphasized that adding the bare pole to the $D^{*}\pi$ amplitude does not affect the narrow pole associated with $D_{1}(2420)$ and only the wider pole changes its position in the complex plane. The ansatz followed in the work, as summarized in Fig 3, is chosen on purpose since a good description of $D_{1}(2420)$ is already obtained with the meson degrees of freedom. The added quark-model pole, on the other hand, in the two models discussed subsequently, becomes a virtual pole  [^6] The pole appears on the second Riemann sheet of the $\bar{K}D_{s}^{*}$, though it lies below the corresponding threshold. and does not leave any trace on the real axis in one of the cases.
@@ II.3.1 Model A
The narrow pole, shown in Fig. 1, remains almost unchanged and the bare quark-model pole becomes a virtual one. The wider pole, obtained at $E-i\Gamma/2=2268-i100$, coincides with the lowest $D_{1}$ state found in unitarized chiral perturbation calculations Du et al. (2018) where the free parameters of the next-leading-order term have been fixed by using lattice QCD calculations. The results of Refs. Abreu et al. (2019); Di Pierro and Eichten (2001) also agree with Ref. Du et al. (2018). Let us discuss the effect of the choice of parameters, together with the negative sign, shown in Eq. (21), on the real axis. For this, we show the squared amplitudes for the different channels in Fig. 4.
@@
As can be seen in the left panel of the top row of Fig. 4, the $D^{*}\pi$ amplitude shows a peak on the real axis around 2304 MeV, with a full width at half maximum of around 160 MeV. Such a width is more in agreement with the lower limit determined by the Babar Collaboration Aubert et al. (2006). The same amplitude shows a zero near 2400 MeV, which is an effect of the negative interference between the lower energy pole arising from meson-meson dynamics and the quark-model pole. Further, it can be seen that the $D\rho$ amplitude is almost the same in Fig. 2 and 4. Thus, clearly, the $D^{*}\pi$ amplitude is dominated by the wider $D_{1}$ state and does not show any clear sign of $D_{1}(2420)$, while the $D\rho$ amplitude shows only the presence of the narrow pole. In fact we find that the $D^{*}\pi$ amplitude gets very little contribution from the coupled channel interactions. Some of the transition amplitudes, like $D^{*}\pi\to D\rho$, do show the presence of an interference effect between a bump and a narrow peak. Such findings are in consonance with the couplings found for the two $D_{1}$ states to the different channels, as listed in Table 2. These couplings have been determined by calculating the residues of the $t$-matrices in the complex energy plane.
We find it relevant to provide the values of the scattering length for different channels also (see Table 2). As discussed earlier, the value for $D^{*}\pi$ is especially interesting since it can be related to the information available on the $D\pi$ channel. We find a value that is in line with the $D\pi$ scattering lengths determined by effective theories [given in Eq. (19)].
Table 2: Values of the isospin 1/2 scattering lengths and the couplings, represented as $g$, of the two states $D_{1}$ for the different channels.
$a^{(1/2)}$ $g_{D_{1}(2430)}$ $g_{D_{1}(2420)}$
(fm) (MeV) (MeV)
$D^{*}\pi$ $-0.20$ $49957250-i$ 4995 $-233+i5$
$D\rho$ $0.180.44-i$ 0.18 $-521-i355$ $35615144+i$ 356
$\bar{K}D_{s}^{*}$ $0.120.00-i$ 0.12 $36124534-i$ 3612 $-247-i177$
$D_{s}\bar{K}^{*}$ $0.120.00-i$ 0.12 $552-i$ 55 $-8739+i90$
@@ II.3.2 Model B
Contrary to Model A, where the mass of a bare pole is taken from a quark model and the coupling, as well as the sign, are chosen so as to determine a pole consistent with Refs. Du et al. (2018); Abreu et al. (2019); Di Pierro and Eichten (2001), we now consider the parameters of Eq. (16) as free. We allow them to vary such as to move the lower energy pole shown in Fig. 1 deeper in the complex plane, thus, providing the possibility of associating a bigger width to the state to be related to $D_{1}(2430)$. As a result, we obtain the following parameterization
$V_{QM}=\frac{10000^{2}}{s-2370^{2}}.$ (22)
The precise position of the pole related to $D_{1}(2430)$ is found to be $2182218-i$ 218 MeV. In this case, a broad bump is found on the real axis around 2436 MeV with a full width at half maximum being $\sim$ 311 MeV (see Fig. 5).
@@
This covers the possibility of a positive interference between the pole arising from the meson-meson dynamics and the bare quark model pole which becomes a virtual pole when the scattering equation is solved. The resulting mass and width values are more in agreement with those found by the LHCb and Belle Collaborations Aaij et al. (2015); Abe et al. (2004) [as also given in Eq. (1)]. Before continuing with further discussions, we must add here that the values of the parameters in Eq. (22) could be changed a bit which would lead to the mass and width still being in agreement with those of $D_{1}(2430)$. Such changes could provide some uncertainty in the amplitudes, but we take here the values of the parameters in Eq. (22) as an example leading to a scenario compatible with the experimental information.
It can be also noticed that the $D\rho\to D\rho$ as well as $D\rho\to D_{s}\bar{K}^{*}$ amplitudes are almost the same as obtained in model A (see Fig. 4), though the strength of the transition of $D\rho$ to the other two channels has diminished. Besides such changes, a cusp effect is seen near 2607 MeV, especially in $D^{*}\pi\to D^{*}\pi$ as well as $D^{*}\pi\to\bar{K}D_{s}^{*}$, which corresponds to the opening of the $\bar{K}D_{s}^{*}$ channel. We must also mention that the $D^{*}\pi$ amplitude, as mentioned in the discussions of model A, gets little contribution from the coupled channel interactions.
In this case, as shown in Table 3, the scattering length of $D^{*}\pi$ turns out to be more in agreement with the value for $D\pi$ determined by the Alice Collaboration [given in Eq. (20)]. We provide the couplings of the two states to the different channels too in Table 3.
Table 3: Values of the isospin 1/2 scattering lengths and the couplings, represented as $g$, of the two states $D_{1}$ for the different channels.
$a^{(1/2)}$ $g_{D_{1}(2430)}$ $g_{D_{1}(2420)}$
(fm) (MeV) (MeV)
$D^{*}\pi$ $0.1$ $35775199-i$ 3577 $10592-i$ 105
$D\rho$ $0.180.45-i$ 0.18 $-248-i91$ $23214987+i$ 232
$\bar{K}D_{s}^{*}$ $0.120.00-i$ 0.12 $22493806-i$ 2249 $39186-i$ 39
$D_{s}\bar{K}^{*}$ $0.120.00-i$ 0.12 $-34-i36$ $-8668+i151$
@@
Notice that the interactions are weakly repulsive in this isospin configuration and, thus, no states are formed in this case. The scattering lengths in this case are $a_{D^{*}\pi}^{(3/2)}=0.1$ fm and $a_{D\rho}^{(3/2)}=0.2$ fm. The value of the isospin 3/2 scattering length, for the $D^{*}\pi$ channel, is in agreement with the one determined for $D\pi$ in other works [as summarized in Eqs. (19) and (20)]. We remind the reader that the our sign convention (as given in section II.2) is opposite to the one followed in the works leading to the values of Eqs. (19) and (20).
@@ III.1 Formalism
where $\vec{k}$ is the relative momentum in the center of mass (CM) of the pair; $\vec{r}$ is the relative distance between the two particles; and $S_{12}(\vec{r})$ is the normalized source function, $\int d^{3}rS_{12}(\vec{r})=1$, describing the distribution of relative positions of particles with identical velocities as they move in their asymptotic state (for a detailed discussion see for example Ref. Lisa et al. (2005)). As a consequence, the expression above for $C(k)$ encodes information on both the hadron source and the hadron-hadron interactions and is commonly named as Koonin–Pratt equation Koonin (1977); Pratt (1986)  [^7] Eq. (23) is also called by some authors like those from Refs. Kamiya et al. (2022) as Koonin–Pratt–Lednicky–Lyuboshits–Lyuboshits formula due to subsequent contributions Lednicky and Lyuboshits (1981); Lednicky et al. (1998).
In the present work, we employ a source function parametrized as a static Gaussian normalized to unity, i.e.
@@
To connect the CF to the coupled-channel approach described in the previous section, we adopt the framework summarized in Refs. Vidana et al. (2023); Feijoo et al. (2023); Albaladejo et al. (2023), in which the generalized coupled-channel CF for a specific channel $i$ reads
@@
where $w_{j}$ is the weight of the observed channel $j$ (we use $w_{j}=1$); $j_{\nu}(kr)$ is the spherical Bessel function; $E=\sqrt{s}$ is the CM energy; the relative momentum of the channel is $k=\lambda^{1/2}(s,m_{1}^{2},m_{2}^{2})/(2\sqrt{s})$ ($\lambda$ being the Källen function and $m_{1},m_{2}$ the masses of the mesons in the channel $i$); $T_{ji}$ are the elements of the scattering matrix encoding the meson–meson interactions, obtained and analyzed in the previous section; and the $\widetilde{G}_{j}(r;s)$ function is defined as

### v2

@@ II.2 Scattering lengths and related uncertainties
at a pion mass of 266 MeV. As can be seen, the values for the two channels are very similar. Such a finding can be understood by invoking arguments of heavy quark symmetry. It is important to stress here that a sign convention opposite to that given in Eq. (17) is followed in Ref. Mohler et al. (2013). Further, scattering lengths for systems like $D\bar{K}$, $D_{s}\pi$, $D_{s}K$, and $D\pi$ in isospin 3/2 configurations have been obtained on a lattice in Ref. Liu et al. (2013). Using the values of the low energy constants of the chiral Lagrangian fixed from such a study, the scattering length for $D\pi$ in isospin 1/2 is determined. The results from the former works, obtained at the physical pion mass, can be summarized as
@@
Recent results on the $D^{*}\pi$ scattering length are available from an alternative source, from the Alice collaboration Acharya et al. (2024), which seem to be in disagreement with the aforementioned values. Let us denote the values given in Refs. Acharya et al. (2024) as $\tilde{a}^{(I)}$. Using such a notation, the values in Refs. Acharya et al. (2024) can be summarized as
$\tilde{a}^{(1/2)}_{D\pi}$ $=0.02\pm 0.03\pm 0.01\text{ fm},$
$\tilde{a}^{(3/2)}_{D\pi}$ $=0.01\pm 0.02\pm 0.01\text{ fm},$
$\tilde{a}^{(1/2)}_{D^{*}\pi}$ $=-0.03\pm 0.05\pm 0.02\text{ fm},$
$\tilde{a}^{(3/2)}_{D^{*}\pi}$ $=0.05\pm 0.04\pm 0.02\text{ fm}.$ (20)
In such a scenario, with different values of scattering lengths obtained from different sources, with one relying on lattice QCD calculations and the other being the experimental data obtained by the Alice collaboration, we find it instructive to investigate if femtoscopic correlation functions measured for source sizes different to the one in Ref. Acharya et al. (2024) can resolve this puzzle. We will eventually show that the correlation functions of $D^{*0}\pi^{+}$ and $D^{*+}\pi^{0}$ are sensitive to the different values of the scattering lengths, and the difference is more marked for smaller source sizes. We hope that our findings can motivate the determination of data satisfying such conditions.
@@ II.3 Amplitudes: Results and discussions
As mentioned in the preceding discussions, we add Eq. (16) to the amplitudes already considered in Refs. Gamermann and Oset (2007); Malabarba et al. (2023) with the idea of having the presence of a wide as well as a narrow $D_{1}$ state around 2430 MeV. In addition for the amplitudes to show states with properties compatible with those of $D_{1}(2420)$ and $D_{1}(2430)$, we require the values of the $D^{*}\pi$ scattering lengths to be in agreement with Eqs. (19) or (20). For this purpose, we shall consider two different sets of values for the parameters $M_{QM}$ and $g_{QM}$, and refer to the cases as models A and B. It must be clarified here that there is little room for changing the only other possible parameter of the model, which is the subtraction constant used to regularize the loop function since the properties of $D_{1}(2420)$ are well described by our amplitudes as discussed in Ref. Malabarba et al. (2023). Note that the scale $\mu$ and the subtraction constant $a_{k}$ appearing in Eq. (15) are not two parameters. Both are related to each other and, hence, together they account for one parameter only.
It should be emphasized that adding the bare pole to the $D^{*}\pi$ amplitude does not affect the narrow pole associated with $D_{1}(2420)$ and only the wider pole changes its position in the complex plane. The ansatz followed in the work, as summarized in Fig 3, is chosen on purpose since a good description of $D_{1}(2420)$ is already obtained with the meson degrees of freedom. Thus, we end up finding two poles which can be related to $D_{1}(2420)$ and $D_{1}(2430)$. The added quark-model pole, on the other hand, in the two models discussed in the following sections, becomes a virtual pole  [^7] A pole which appears below the threshold of a given channel, on the unphysical Riemann sheet, is called as a virtual pole..
@@ II.3.1 Model A
The narrow pole, shown in Fig. 1, remains almost unchanged and the bare quark-model pole becomes a virtual one (appearing at $02448-i$ 0 MeV). The wider pole, obtained at $E-i\Gamma/2=2268-i100$, coincides with the lowest $D_{1}$ state found in unitarized chiral perturbation calculations Du et al. (2018) where the free parameters of the next-leading-order term have been fixed by using lattice QCD calculations. The results of Refs. Abreu et al. (2019); Di Pierro and Eichten (2001) also agree with Ref. Du et al. (2018). Let us show the squared amplitudes for the different channels in Fig. 4 for further discussions.
@@
As can be seen in the left panel of the top row of Fig. 4, the $D^{*}\pi$ amplitude shows a peak on the real axis around 2305 MeV, with a full width at a half maximum of around 160 MeV. Such a width is more in agreement with the lower limit determined by the BABAR collaboration Aubert et al. (2006). The same amplitude shows a zero near 2400 MeV, which is an effect of the interference between the lower energy pole arising from meson-meson dynamics and the quark-model (virtual) pole. Further, it can be seen that the $D\rho$ amplitude is almost the same in Figs. 2 and 4. Thus, clearly, the $D^{*}\pi$ amplitude is dominated by the wider $D_{1}$ state and does not show any clear sign of $D_{1}(2420)$, while the $D\rho$ amplitude shows only the presence of the narrow pole. In fact, we find that the $D^{*}\pi$ amplitude gets very little contribution from the coupled channel interactions. Some of the transition amplitudes, like $D^{*}\pi\to D\rho$, do show the presence of an interference effect between a bump and a narrow peak. Such findings are in consonance with the couplings found for the two $D_{1}$ states to the different channels, as listed in Table 2. These couplings have been determined by calculating the residues of the $t$-matrices in the complex energy plane. For the sake of completeness, we also present the position and couplings of the virtual pole in Table 2.
We find it relevant to provide the values of the scattering length for different channels as well (see Table 2). As discussed earlier, the value for $D^{*}\pi$ is especially interesting since it can be compared with those given in Eqs. (19) and (20).
Table 2: Values of the isospin 1/2 scattering lengths and the couplings, represented as $g$, of the two states $D_{1}$ for the different channels, as obtained in model A. Further, we show the couplings of the virtual pole found in the work (as discussed in the text). The asterisks on $g$ indicate that the values of the pole (in the subscript) and the couplings (in the last column) correspond to a virtual pole.
$a^{(1/2)}$ $g_{D_{1}(2430)}$ $g_{D_{1}(2420)}$ $g^{**}_{2448-i0}$
(fm) (MeV) (MeV) (MeV)
$D^{*}\pi$ $-0.20$ $49957250-i$ 4995 $-233+i5$ $6826302-i$ 6826
$D\rho$ $0.180.44-i$ 0.18 $-521-i355$ $35615144+i$ 356 $-284+i719$
$\bar{K}D_{s}^{*}$ $0.120.00-i$ 0.12 $36124534-i$ 3612 $-247-i177$ $4231346-i$ 4231
$D_{s}\bar{K}^{*}$ $0.120.00-i$ 0.12 $552-i$ 55 $-8739+i90$ $11041-i$ 110
At this point, we find it useful to make a brief discussion on how the couplings (for physical poles) listed in Table 2 can be used in determining observables. For instance, these couplings can lead to partial decay widths, through
$\Gamma_{D_{1}\to m_{1}m_{2}}=\frac{p_{c.m.}}{8\pi M_{D_{1}}^{2}}\left|g_{D_{1}}\right|^{2},$ (22)
[…]
To obtain more precise values, we must take into account that $M_{D_{1}}$ varies in the range allowed by the related finite width of the decaying particle (as shown in Ref. Oller and Oset (1997)) and calculate the partial decay width as
[…]
@@ II.3.2 Model B
Contrary to Model A, where the mass of a bare pole is taken from a quark model and the coupling, as well as the sign, are chosen so as to determine a pole consistent with Refs. Du et al. (2018); Abreu et al. (2019); Di Pierro and Eichten (2001), we now consider the parameters of Eq. (16) as free. We allow them to vary such as to move the lower energy pole shown in Fig. 1 deeper in the complex plane, thus, providing the possibility of associating a bigger width to the state to be related to $D_{1}(2430)$. As a result, we obtain the following parametrization
$V_{QM}=-\frac{10000^{2}}{s-2370^{2}}.$ (24)
The precise position of the pole related to $D_{1}(2430)$ is found to be $2182281-i$ 218 MeV. In this case, a broad bump is found on the real axis around 2436 MeV with a full width at half maximum being $\sim$ 311 MeV (see Fig. 5).
@@
In this case, there occurs a constructive interference between the pole arising from the meson-meson dynamics and the bare quark model pole which becomes a virtual pole when the scattering equation is solved. The resulting mass and width values, determined from the peak on the real axis, are more in agreement with those found by the LHCb and Belle collaborations Aaij et al. (2015); Abe et al. (2004) [as also given in Eq. (1)].
It can also be noticed that the $D\rho\to D\rho$ as well as $D\rho\to D_{s}\bar{K}^{*}$ amplitudes are almost the same as obtained in model A (see Fig. 4), though the strength of the transition of $D\rho$ to the other two channels has diminished. Besides such changes, a cusp effect is seen near 2607 MeV, especially in $D^{*}\pi\to D^{*}\pi$ as well as $D^{*}\pi\to\bar{K}D_{s}^{*}$, which corresponds to the opening of the $\bar{K}D_{s}^{*}$ channel. We must also mention that the $D^{*}\pi$ amplitude, as mentioned in the discussions of model A, gets little contribution from the coupled channel interactions.
In this case, as shown in Table 3, the scattering length of $D^{*}\pi$ turns out to be more in agreement with the value determined by the Alice collaboration [given in Eq. (20)]. We provide the couplings of the two states to the different channels, as well as to the virtual pole, too in Table 3.
Table 3: Values of the isospin 1/2 scattering lengths and the couplings, represented as $g$, of the two states $D_{1}$ and of the virtual pole (marked by asterisks) for the different channels, as obtained in model B.
$a^{(1/2)}$ $g_{D_{1}(2430)}$ $g_{D_{1}(2420)}$ $g^{**}_{2370-i0}$
(fm) (MeV) (MeV) (MeV)
$D^{*}\pi$ $0.1$ $35775199-i$ 3577 $10592-i$ 105 $-1489-i5244$
$D\rho$ $0.180.45-i$ 0.18 $-248-i91$ $23214987+i$ 232 $-182+i870$
$\bar{K}D_{s}^{*}$ $0.120.00-i$ 0.12 $22493806-i$ 2249 $39186-i$ 39 $932945-i$ 9329
$D_{s}\bar{K}^{*}$ $0.120.00-i$ 0.12 $-34-i36$ $-8668+i151$ $20740-i$ 207
@@
Notice that the interactions are weakly repulsive in this isospin configuration and, thus, no states are formed in this case. The scattering lengths in this case are $a_{D^{*}\pi}^{(3/2)}=0.1$ fm and $a_{D\rho}^{(3/2)}=0.2$ fm. The value of the isospin 3/2 scattering length, for the $D^{*}\pi$ channel, is in agreement with the ones summarized in Eqs. (19) and (20). We remind the reader that our sign convention (as given in Sec. II.2) is opposite to the one followed in the works leading to the values of Eqs. (19) and (20).
@@ III Correlation Functions
As stated earlier the purpose of our work is to determine the correlation function for the $D^{*}\pi$ system. We focus on investigating the $D^{*+(0)}\pi^{0(+)}$ system where Coulomb interactions are absent and strong interactions dominate. The idea is also to study the dependence of the correlation function on the size of the source, such as to find if experimental data on processes other than that studied in Ref. Acharya et al. (2024) can bring useful information on the topic.
@@ III.1 Formalism
where $\vec{k}$ is the relative momentum in the c.m. of the pair; $\vec{r}$ is the relative distance between the two particles; and $S_{12}(\vec{r})$ is the normalized source function, $\int d^{3}rS_{12}(\vec{r})=1$, describing the distribution of relative positions of particles with identical velocities as they move in their asymptotic state (for a detailed discussion see for example Ref. Lisa et al. (2005)). As a consequence, the expression above for $C(k)$ encodes information on both the hadron source and the hadron-hadron interactions and is commonly named as Koonin–Pratt equation Koonin (1977); Pratt (1986)  [^8] Eq. (25) is also called by some authors like those from Refs. Kamiya et al. (2022) as Koonin–Pratt–Lednicky–Lyuboshits–Lyuboshits formula due to subsequent contributions Lednicky and Lyuboshits (1981); Lednicky et al. (1998).
In the present work, we employ a source function parametrized as a static Gaussian normalized to unity, i.e.,
@@
To connect the CF to the coupled-channel approach described in the previous section, we adopt the framework summarized in Refs. Vidana et al. (2023); Feijoo et al. (2023); Albaladejo et al. (2023); Liu et al. (2023a, b, c), in which the generalized coupled-channel CF for a specific channel $i$ reads
@@
where $w_{j}$ is the weight of the observed channel $j$ (we use $w_{j}=1$); $j_{\nu}(kr)$ is the spherical Bessel function; $E=\sqrt{s}$ is the c.m. energy; the relative momentum of the channel is $k=\lambda^{1/2}(s,m_{1}^{2},m_{2}^{2})/(2\sqrt{s})$ ($\lambda$ being the Källen function and $m_{1},m_{2}$ the masses of the mesons in the channel $i$); $T_{ji}$ are the elements of the scattering matrix encoding the meson–meson interactions, obtained and analyzed in the previous section; and the $\widetilde{G}_{j}(r;s)$ function is defined as